OpenStudy (anonymous):

Find the general solution of the differential equation \[{dy\over dx} = {{3x^2 + y}\over{2y^2 + x}}\]

OpenStudy (turingtest):

this is going to be one of those things where we divide top and bottom by x and sub in v=y/x I think...

OpenStudy (anonymous):

okay, I'll give it a go

OpenStudy (turingtest):

I'm not so good at these, but here is a good set of notes on this stuff http://tutorial.math.lamar.edu/Classes/DE/Substitutions.aspx

OpenStudy (amistre64):

i wonder if we can exact it

OpenStudy (turingtest):

yeah I'm getting stuck with my sub it looks exact

OpenStudy (amistre64):

(2y^2+x) dy - (3x^2+y)dx=0

OpenStudy (anonymous):

I think that it would be exact if it wasn't for the minus sign

OpenStudy (anonymous):

I got stuck with that substitution too

OpenStudy (amistre64):

yeah, when i do exact i keep kitting a spurious 2 in there

OpenStudy (turingtest):

\[v=y/x\]\[v+xv'={3x+v\over2xv^2+1}\]\[2xv^3+2x^2v^2v'+v+xv'=3x+v\]I don't suppose this can be separated or linearized...

OpenStudy (anonymous):

Thats where I got stuck with that substitution, I can't see what to do with that

OpenStudy (turingtest):

wolfram isn't helping either http://www.wolframalpha.com/input/?i=solve+dy%2Fdx%3D%283x%5E2%2By%29%2F%282y%5E2%2Bx%29

OpenStudy (turingtest):

all I can do is request backup @lalaly @myininaya DE help?

OpenStudy (amistre64):

yeah, the wolf gives up too; we can try the euler method :)

OpenStudy (amistre64):

2y(x)^2 + x = 0 y(x)^2 = -x/2 y(x) = sqrt(-x/2) ; looks like its good across the reals at least

OpenStudy (amistre64):

draw a slope field?

OpenStudy (anonymous):

I don't think that will cut it, we haven't even covered that in this class :)

OpenStudy (amistre64):

slope fields is like the first thing covered in class

OpenStudy (anonymous):

I remember it being mentioned in calculus 2 but not everyone has seen it in this DE class

OpenStudy (amistre64):

you sure you typed it up correctly?

OpenStudy (anonymous):

Yes, I've checked and rechecked. I've made that mistake before

OpenStudy (amistre64):

\[\frac{dy}{dx}=\sum_{n=0}^{inf}3(-2)^nx^{1-n}y^n\] not that it helps ...

OpenStudy (amistre64):

and i mighta messed up the long division

OpenStudy (anonymous):

- (3x^2+y)dx + (2y^2+x) dy=0 it is exact ODE

OpenStudy (anonymous):

1º: integrate - (3x^2+y)dx on x 2º diferentiate the result by y 3º make equal to (2y^2+x) to find the constant and good job

OpenStudy (amistre64):

how do you get over that -1 not= 1?

OpenStudy (anonymous):

.....:(

OpenStudy (anonymous):

See I am not sure whether this is the right method or not, but please check \[\int\limits_{}^{}(2y^2+x)dy=\int\limits_{}^{}(3x^2+y)dx\] \[2\int\limits_{}^{}y^2dy+x \int\limits_{}^{}dy= 3\int\limits_{}^{}x^2dx +y \int\limits_{}^{}dx\] \[2{{y^3}\over{3}} + xy = 3{{x^3}\over{3}} +xy+C\] \[2{{y^3}\over{3}} = 3{{x^3}\over{3}}+c \]

OpenStudy (amistre64):

tried it

OpenStudy (anonymous):

Is it right that I can assume x to be constant while integrating with dy ??

OpenStudy (amistre64):

i think you need to account, not for a C but for a f(x) for the dy and an g(y) for the dx

OpenStudy (amistre64):

you can, but the resulting integrations dont necessarily have an aribitrary C, the have an arbitrary function of x or y that zeros out instead

OpenStudy (anonymous):

But it's not exact, you can't just integrate both sides and find f(x) and g(y)

OpenStudy (amistre64):

but since the Fxy and Fyx arent exactly the same, it throws in a monkey wrench

OpenStudy (anonymous):

@amistre64 , thanks for confirmation :) I will try this sum in a different method

OpenStudy (turingtest):

where did this problem come from?

OpenStudy (anonymous):

I've been wondering if there is some integrating factor that would make this exact but haven't had any luck finding one

OpenStudy (amistre64):

if you ignore the "-" can we get it to work??

OpenStudy (anonymous):

This is the take-home portion of our midterm. We are encouraged to work on them in groups

OpenStudy (turingtest):

@amistre64 I'm pretty sure we can't ignore the -

OpenStudy (anonymous):

This thing would be so trivial if it wasn't for that darn -

OpenStudy (amistre64):

\[C_0=-x^3-xy+xy+\frac{2}{3}y^3+C_1\] seems to work

OpenStudy (turingtest):

where did that come from o-0 ?

OpenStudy (amistre64):

lol, playing around with some ideas; its prolly wrong tho

OpenStudy (amistre64):

i dont see a way to get to it ....

OpenStudy (amistre64):

you got any initial conditions?

OpenStudy (anonymous):

no, no initial conditions

OpenStudy (amistre64):

you can get back at your teacher by asking them to solve it :)

OpenStudy (anonymous):

I could definitely use some kind of hint! Hopefully I can get one from her :)

OpenStudy (anonymous):

@satellite73 , Can you help please??

OpenStudy (anonymous):

maybe this helps: opening parentesis: \[-3x ^{2}dx -ydx +2y ^{2}dy +xdy=0\] grouping terms trying to get exact differentials: \[[-ydx+xdy] -[3x ^{2}dx-2y ^{2}dy] =0\] the first term is equal to : \[x ^{2} d(y/x)\] i bet 2º one is something similar...

OpenStudy (anonymous):

@myko , The second one cannot be written as you have written in first one or anything similar.

OpenStudy (anonymous):

i think it can

OpenStudy (anonymous):

See the position of dx and dy in first one and second one.

OpenStudy (anonymous):

it should be the differential of oposit: d(x/y), but maybe squared or somthing like that

OpenStudy (anonymous):

i saw this type of equations befor. You solve them this way....

OpenStudy (anonymous):

Not possible Even of form d(xy) , d(x/y) . Nothing can be possible. Had it been 3x^2 dy−2y^2dx then the story would have been different but it is 3x^2dx−2y^2dy :(

OpenStudy (anonymous):

just muliply the diffrential by x^2/y^2 and you got it

OpenStudy (anonymous):

??

OpenStudy (anonymous):

like i did with first one...

OpenStudy (anonymous):

I just got an email back from the instructor and there was a typo, it should have been: \[{dy \over dx} = {{3x^2 + y \over 2y^2 - x}}\] It is exact!

OpenStudy (lalaly):

lol so u still need help with it or u got it?